Find the midpoint of the line segment with the given endpoints: (2, 4) and (1, -3).
1
Expert's answer
2012-09-13T08:43:26-0400
xm=(2+1)/2=1.5
ym=(4-3)/2=0.5
So the midpont is M(1.5, 0.5)